OpenStudy (anonymous):

The solution to the radical function √x-6=2 is x = 8. True False

OpenStudy (anonymous):

False, √x=2+6 x = 8^2 x = 64

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Your welcome :D